OpenStudy (anonymous):

x^2-3x+5=x(x-8)+20 help!

OpenStudy (anonymous):

\[x^2-3x+5=x^2-8x+10\] \[-3x+5=-8x+10\] is a start

OpenStudy (anonymous):

x=3

OpenStudy (lgbasallote):

i think @satellite73 means -3x + 5 = -8x + 20 :P

OpenStudy (anonymous):

so the answer is x=3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did u get that answer?

OpenStudy (anonymous):

Just as @satellite73 and @latineclipse depicted their equations, then solve/factor for "x"

OpenStudy (anonymous):

and u got 3 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thnx so much